can you help solve: 2/5x-3y=8; 4x+y=18 for (x,y)?
Answer by stanbon(75887)

solve: (2/5)x-3y=8; 4x+y=18 for (x,y)
---------------------
Rearrange:
2x - 15y = 50
4x + y = 18
----------------
Modify:
4x - 30y = 100
4x + y = 18
---------------
Subtract 1st from 2nd to get:
31y = -82
y = -82/31
-------------
Substitute into 4x+y = 18 to solve for "x":
4x - (82/31) = 18
4x = 18 + (82/31)
x = 160/31
----------------
Cheers,
Stan H.